Two objects attract each other with a gravitational force of magnitude 1.00 10-8 N when separated by 19.3 cm. If the total mass of the objects is 5.05 kg, what is the mass of each?


1
Expert's answer
2021-10-31T18:15:20-0400

"F = 1.0 \times 10^{-8} \;N \\

r = 19.3 \times 10^{-2} \;m \\

m_1+m_2=5.05"

From Newton’s law of force

"F = \frac{Gm_1m_2}{r^2} \\

1.0 \times 10^{-8} = \frac{6.67 \times 10^{-11} \times m_1m_2}{(19.3 \times 10^{-2})^2} \\

5.584 = m_1m_2 \\

m_1+m_2=5.05 \\

m_1 = 5.05 -m_2 \\

5.584 = (5.05 -m_2)m_2 \\

5.584 = 5.05m_2 -m_2^2 \\

m^2_2 -5.05m_2 + 5.584 = 0 \\

m_2 = 3.41 \\

m_1 = 1.64"